OpenStudy (mandre):

Area between graphs. I thought this would be simple until I tried to calculate it. The area between f(x) = x^2 - 6 and g(x) = |x|

OpenStudy (phi):

What have you tried ?

OpenStudy (tkhunny):

Did you solve your problem with the absolute values by observing symmetry about the y-axis?

OpenStudy (phi):

Here is a graph of the two curves

OpenStudy (mandre):

I tried calculating each area (between graph and x-axis) and subtracting but I don't think that will work

OpenStudy (mandre):

Correction. it should be -|x|

OpenStudy (phi):

if we look at the right half for x≥0 the area of a thin rectangle will have height g(x)- f(x) or x-( x^2 - 6) and width dx do you know how to integrate from x=0 to 3 this expression ?

OpenStudy (phi):

if g(x) = -|x| then the graph is this

OpenStudy (phi):

and the height of each rectangle is -x - (x^2 -6) notice that the integration should go from 0 to x= 2

OpenStudy (phi):

and then double the area of the right side for the total area.

OpenStudy (mandre):

I tried Wolfram Alpha but they don't show how to do it. \[\int\limits_{-2}^{2} x ^{2}-6-|x| dx = \int\limits_{-2}^{0} x ^{2}-6-(-x) dx + \int\limits_{0}^{2} x ^{2}-6-(x) dx \] Is that right? Then I work out one half and double?

OpenStudy (phi):

yes, you have to split the integral because |x| is really two different functions depending on what x is. So you can do what you posted, or integrate one or the other and double the result, because the two parts have identical area (by symmetry) but I would make the integral for x=0 to 2: -x - (x^2 -6) = -x^2 -x + 6

OpenStudy (mandre):

I get 22/3 which will give me 44/3 and that is one of my options. Multiple choice.

OpenStudy (phi):

yes. The only detail is whether we should make the area negative. Area below the x-axis is often made negative... on the other hand, if all the choices area positive, then go with +44/3